OpenStudy (anonymous):

The number of internet users I in millions during year x, where x > 2000 can be approximated by the following formula I= 118x- 235,529 Approximate the year in which there were 600 million internet users for the first time. I came up with the year 1996 and it was incorrect for some reason

OpenStudy (anonymous):

no, I did it like this 600= 118x - 235,529 235,411=118x 235,411/118 = 118x/118 235,411/118= x x = 1995

OpenStudy (cwrw238):

oh - i didn't read question through sorry

OpenStudy (cwrw238):

how did you get 235411 in 2nd line

OpenStudy (anonymous):

I subtracted 118 from 235,529

OpenStudy (cwrw238):

no can't do that you have a number 235529 and term in 118x - you can't add/subtract unlike terms

OpenStudy (anonymous):

ok

OpenStudy (cwrw238):

600= 118x - 235,529 add 235529 to both sides of the '=' sign: 236129 = 118x x = 236129 / 118

OpenStudy (anonymous):

oh ok so the answer would be 2001, right?

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

thanks, this algebra is giving me a nervous break down, haha

OpenStudy (cwrw238):

yea - its important to remember you can only add like terms eg x + 12x = 13x 2 + 12x = 2 + 12x: can't simplify that any further

OpenStudy (cwrw238):

practice as much as you can - your confidence will increase

OpenStudy (anonymous):

I just dont know how people remember all these rules. I think I going to have to make me a notebook of all the rules and set down and memorize them , lol

OpenStudy (cwrw238):

yes - that will help